OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6).

OpenStudy (anonymous):

just plug in f(x) = -2x + 1 f(-6) = -2(-6) + 1 = 12 + 1 =13

OpenStudy (anonymous):

just plug in -6 for x. so: -2(-6) + 1 can you solve from there?

OpenStudy (anonymous):

no >.< im so confused when it comes to algebra

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

its just add and subtracting numbers